OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=4xsin(x) at the point (π/2,2π). Express your answer in the form y=mx+b. i know the derivative is 4sinx+4xcosx

hartnn (hartnn):

derivative at point \((\pi/2 , 2\pi) \) will give you the slope of tangent at that point

hartnn (hartnn):

so plug in x = pi/2 in your derivative

OpenStudy (anonymous):

yes but i got the wrong answer

hartnn (hartnn):

what slope did u get ?

OpenStudy (anonymous):

4

hartnn (hartnn):

thats correct y - 2pi = 4 (x - pi/2) you got this ? simplified ? what did u get ?

OpenStudy (anonymous):

y=4(x-pi/2) +2pi

hartnn (hartnn):

yep multiply 2 on both sides

hartnn (hartnn):

or distribute 4 to (x-pi/2)

hartnn (hartnn):

do u get the final answer as y =4x ?

OpenStudy (anonymous):

Pi/2 then

OpenStudy (anonymous):

yes i got the answer

hartnn (hartnn):

cool, any more doubts ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^